OpenStudy (anonymous):

sec theta = -4√7/7 Find sin(2theta) if cdc theta is greater than 0

OpenStudy (misty1212):

hi

OpenStudy (misty1212):

draw a triangle, make the hypotenuse \(4\sqrt7\) and the adjacent side \(7\)

OpenStudy (misty1212):



OpenStudy (misty1212):

all that is missing is \(b\) which you find by pythagoras \[b=\sqrt{(4\sqrt{7})^2-7^2}=\sqrt{63}\]

OpenStudy (misty1212):



OpenStudy (misty1212):

now you can take any trig ratio you want

OpenStudy (misty1212):

ooh you want \[\sin(2\theta)\]

OpenStudy (misty1212):

that is \[2\sin(\theta)\cos(\theta)\] we can find it in a second

OpenStudy (anonymous):

Thank you!

OpenStudy (misty1212):

but first what does cdc theta is greater than 0 mean?

OpenStudy (misty1212):

anyways \[\cos(\theta)=-\frac{7}{4\sqrt7}\] and \[\sin(\theta)=\frac{\sqrt{63}}{4\sqrt{7}}\]

OpenStudy (misty1212):

plug them in above and see what you get when you simplify \[2\times -\frac{7}{4\sqrt7}\times \frac{\sqrt{63}}{4\sqrt{7}}\]

OpenStudy (anonymous):

how did you get sqrt 63?